[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Sep 20, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Sep 20 19:37:14 UTC 2016


HSU Shift Log for Tue Sep 20 2016 15:37:08 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues: Pacific desk ATCF workstation was having some screen refresh  
issues when windows were being rendered. Stacy rebooted the workstation and  
it seemed to be better after that.

N-AWIPS/AWIPS2 issues: None

Other Techincal issues: Several issues from last night and early this  
morning were related to problems with the naprod account on one of the  
compute servers (comp01), including ASCAT-B data ingest, ingest of UKMET  
model grids for NAWIPS, and ingest of HCCA into the ATCF. TSB is aware of  
the issue and has switched to the other compute server (comp03) for now.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160920/8ac02548/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list